OpenStudy (anonymous):

Need help with 15 8th grade questions. Will medal and fan!

OpenStudy (anonymous):

ill do my best to help you out :D

OpenStudy (anonymous):

What is the solution to the equation?

OpenStudy (anonymous):

-18+4+x=9-19 x=

OpenStudy (anonymous):

so keep x on one side and move all the the other number to the other... you should end up with x=9-19+18-4, add them all up and there you have it :D

OpenStudy (anonymous):

So, 4?

OpenStudy (anonymous):

@Emeyluv99

OpenStudy (anonymous):

4x-26-3x=16-20 x=

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

now for this second one, keep all the x's on one side and move everything else to the other

OpenStudy (anonymous):

4-26-3=16-20? @emeyluv99

OpenStudy (anonymous):

hang on, where are your x's you should be getting 4x-3x= 16-20+26

OpenStudy (anonymous):

I don't know. I'm confused :x I don't understand this.

OpenStudy (anonymous):

ok so you want to figure out the value of x. Yes?

OpenStudy (anonymous):

Correctumundo.

OpenStudy (anonymous):

so you have two numbers that are multiples of x, but, because you dont know what x is, you dont know what 4x or 3x is.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so now, what we can do is put all the numbers that we know the value of (ie all the number that are not multiples of x) on one side, and keep the multiples of x on the other

OpenStudy (anonymous):

this just helps us see what we do and dont know and makes it easier to solve.

OpenStudy (anonymous):

go ahead and try it.

OpenStudy (anonymous):

So, 4x-3x= 16-20+26

OpenStudy (anonymous):

@Emeyluv99

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

I just copied what you put earlier....xd

OpenStudy (anonymous):

so now, what do you think 4x-3x is

OpenStudy (anonymous):

(ahah cheeky)

OpenStudy (anonymous):

I don't know!~ I am SO confused. No matter how it gets explained I don't understand.

OpenStudy (anonymous):

ok lets try some experimenting then shall we?

OpenStudy (anonymous):

We shall.

OpenStudy (anonymous):

so lets pretend that x=5.

OpenStudy (anonymous):

whats 3*5-4*5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

NO!

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yes, youre correct

OpenStudy (anonymous):

its - 5

OpenStudy (anonymous):

15-20= -5

OpenStudy (anonymous):

now try with x=20. whats 4*20-3*20

OpenStudy (anonymous):

20

OpenStudy (anonymous):

good

OpenStudy (anonymous):

how about when x=30? whats 4*30-3*30?

OpenStudy (anonymous):

120-90=30

OpenStudy (anonymous):

I get this part.

OpenStudy (anonymous):

ae you seeing a pattern here?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so theres this rule in algebra. It goes somehting like this: ax+bx =(a+b)x

OpenStudy (anonymous):

here a can be any number and b can be any number. id either of them are negative, that meass youre subtracting.

OpenStudy (anonymous):

in your case, a is 4 and b is -3. see if you can work out what 4x-3x is with this rule

OpenStudy (anonymous):